Project

General

Profile

2018-06-12-MEMOMIPO » History » Version 6

Rogers, Chris, 12 June 2018 16:11

1 1 Long, Kenneth
h1. 2018-06-12-MEMOMIPO
2
3
h2. Joint MICE Experiment Management Office and MICE International Project Office
4
5
h3. 12th June 2018; 15:30 BST
6
7
"Phone details":http://mice.iit.edu/phonebridge.html
8
9
h4. Agenda
10
11
# *Introduction: KL*
12
# *Notes of previous meeting and status of actions: All*
13
** *PS:* Generate two-pager to close-out common fund for next PSAG.
14
** *DR:* Continue to push towards a persistence plan for the MICE IIT WWW site.
15
** *DR:* Coordinate migration of mail server from IIT to jiscmail.
16
# *Project Manager's report: CW*
17
# *Software and computing status: DR*
18
# *Paper status and plans: CR*
19
# *AOB*
20
21 2 Rogers, Chris
h4. Notes
22
23
Present:
24
Chris Rogers
25
Peter Garbinicius
26
Colin Whyte
27
Durga Rajaram
28
29 6 Rogers, Chris
# *Introduction - KL* is in another meeting, CR to chair
30 2 Rogers, Chris
# *Notes of previous meeting and status of actions: All*
31
** *PS:* Generate two-pager to close-out common fund for next PSAG.
32
*** No report - stands
33
** *DR:* Continue to push towards a persistence plan for the MICE IIT WWW site.
34
*** Purchase order has been placed for webserver at IIT. Static mirror at RAL to be coordinated with Chris Brew.
35
** *DR:* Coordinate migration of mail server from IIT to jiscmail.
36
*** KL will migrate EB and CB mailing lists to jiscmail
37
*** CR will migrate analysis mailing list to jiscmail
38
*** DR will migrate software mailing list to jiscmail
39
*** Other mailing lists are inactive and will be archived
40
# *Project Manager's report: CW*
41
** See slides
42
** Since VC
43
*** Tony will go into the synchrotron at next opportunity to check a few survey things
44
*** ISIS have offered resource to help clear MICE hall
45
*** Shipping container for US equipment now expected to arrive w/c 25th June; Alan Br likely to be available in that week
46
# *Software and computing status: DR*
47
** V.Pec has generated a set of new TOF calibrations. TOF bad channel patch has also been brough in. DR running some tests, in time for next release.
48
** 2017/03 geometry has wrong alignment corrections; fix is in hand.
49
** Global track matching for backtracking from TKD is now in test. Will not make next release.
50 3 Rogers, Chris
** Discussion regarding "cuts" routines; they should be run by default; DR asked CR to check the default values are reasonable
51
** Handover of computing project
52
*** Archiving mostly done
53
*** Needs some one to manage any data issues and new users (for GRID)
54 5 Rogers, Chris
*** Durga needs to hand over geometry and "all those things". Some discussion of the details. *Action: DR* to make a list.
55 2 Rogers, Chris
# *Paper status and plans: CR*
56
** No update since VC
57
# *AOB*
58
** No AOB
59
60 1 Long, Kenneth
h4. Summary of actions:
61
62 6 Rogers, Chris
* *PS:* Generate two-pager to close-out common fund for next PSAG.
63
* *DR:* List the tasks that need to be handed over for software and computing.
64 1 Long, Kenneth
h3. Back to MEMO:
65
66
[[memo:|Back]]
67
68
h3. Back to MIPO:
69
70
[[mipo:|Back]]
71
72
h2. Documents